Session: Feb 2006 (session 1)
Wavelength: 18 cm


To give feedback on this experiment, click

Experiment Feedback

Urumqi (Zhang Hua at Sun Mar 5 06:51:37 2006) Minor failures : RFI

Medicina (Fiocchi Franco at Sun Mar 5 11:39:33 2006) Minor failures : WIND HIGH POIN
Pointing problems due high speed wind.

Hartebeesthoek (Jonathan Quick at Sun Mar 5 11:47:26 2006) Minor failures : WIND RFI
Persitant spike at 132.5 MHz in IF especially RCP, otherwise repeated bursts of broadband noise. Wind stowed from 064.04:23-04:28 UT

Onsala (Michael Lindqvist at Sun Mar 5 12:50:46 2006) Success :
RFI seen in the bandpass. Otherwise no known problems.

Torun (Kaz Borkowski at Mon Mar 6 06:33:50 2006) Success :
No known problems.

Effelsberg (A. Polatidis at Mon Mar 6 07:42:53 2006) Success :
Ended early at 0756 UT, due to receiver change.

Westerbork (Ton Foley at Mon Mar 6 13:05:05 2006) Success :
Wb used in phased array mode. OK

Jodrell_Bank (Alastair Gunn at Mon Mar 6 13:08:21 2006) Success :
The Lovell Telescope was used (Jb1). No known problems.

Cambridge (Alastair Gunn at Mon Mar 6 13:09:37 2006) Success :
No known problems.

Robledo (Cristina Garcia Miro at Tue Mar 7 13:19:33 2006) Minor failures : WIND RFI
Very bad weather (rain+wind+snow) but not much impact on L-Band. gps-fmout value at start=+4.326 10-7. VC05 output degraded by L-band RFI. Recording started at 0259UT (official starting time for Ro was 0330UT). Labels: JIVE-041, OSOD-019.

Noto (P. Cassaro at Tue Mar 7 16:20:59 2006) Minor failures : HIGH

EVN experiment feedback facility: entire session

Entire Session

Session: Feb 2006 (session 1)
Wavelengths: 6 3.6 5 18 cm


To give feedback on this entire session, click

Entire Session Feedback

Torun (Kaz Borkowski at Mon Feb 20 11:25:29 2006) Minor failures : PCAL
Tr phase cal system failed to produce signals during all the 6 cm experiments for yet unknown reason.

Westerbork (Tony Foley at Fri Feb 24 15:48:24 2006) Minor failures : ANT POIN
The feedback from the antenna control for SINGLE-DISH observations is unreliable for 20seconds after a source change. We will try to fix this. (Affected only the 5cm experiments -Z.P.)